OpenStudy (bookworm14):

I completely forgot how to do these ugh !!!! can you try to help me? (i will post it below in comments) @satellite73 @tester97 @ganeshie8

OpenStudy (bookworm14):

Help?

OpenStudy (bookworm14):

@tester97 @ganeshie8

OpenStudy (cwrw238):

try and solve them and see what you get

OpenStudy (bookworm14):

i am working on it now

OpenStudy (cwrw238):

if you get something like 2 = 5 then there is no solution

OpenStudy (bookworm14):

should i be trying to solve for the variable ex: 7w=8w, or should i be trying to solve for numbers ex: 1=2

OpenStudy (cwrw238):

solve for the variable

OpenStudy (bookworm14):

ok thats what i was thinking

OpenStudy (cwrw238):

it might be that the variable disappears...

OpenStudy (bookworm14):

could you maybe work out a problem as an example (it doesn't have to be one off the question)

OpenStudy (anonymous):

4a +6=4a-2 6=-2 it has no solution.

OpenStudy (bookworm14):

@cwrw238 @surjithayer I got D. 8y+7 = 8y-5 has no solution... is this correct?

OpenStudy (cwrw238):

yes